# Machine-assisted classification of potential biosignatures in earth-like exoplanets using low signal-to-noise ratio transmission spectra

### v1

@@ 6.3.1 SC multilabel metrics
In Figure 12, we show the result of calculating the multilabel classification metrics for the SC. If we compare these results with those corresponding to the MC (see Figure 7) we notice that, despite the differences between the approaches in terms of complexity of the training set and scalability, the performance is almost the same. This implies that training a set of specialised classificators, with its advantages of adding new molecules as required, produces almost the same results, at least for the test spectra we are generating here, then a more complex and CPU intensive algorithm that includes all molecules at once (the MC algorithm).
Still, in a more general case with a realistic spectrum, the MC approach will show its advantages over the SC as we will see in section 8.
Figure 12: Multilabel metrics for the SC algorithm as a function of SNR. For the meaning of the shaded region, see Figure 3 .
@@ 6.3.2 SC molecule metrics
In Figure 13, we perform a similar comparison but for the metrics evaluated for each molecule separately. The analogous results for the MC were presented in Figure 8. Again the performance of both algorithms is comparable, at least for the biosignature molecules. This implies, again, that when classifying molecules individually, the more scalable SC algorithm could be as good as the MC one.
Figure 13: Molecule-specific metrics from SC. Each subplot show the performance of different metrics (F1 Score, Recall, Precision, TNR) for each molecule (CH4, O3, H2O) as the SNR increases. For the meaning of the shaded region, see Figure 3.
Interestingly, and as an example of the “hidden” behaviour of the algorithm, the TNR metric for CH4 and H2O (lower-right panels) tends to stagnate or worsen slightly after a certain SNR value. This behaviour indicates that as the signal becomes less noisy, the algorithm tends to confuse these molecules with others. This is likely due to those other molecules sharing many spectral peaks with the molecules of interest, leading to false positives. However, O3 does not exhibit this behaviour, allowing it to maintain a consistently improving precision across all SNR values.
@@ 6.3.3 SC minimum mixing ratio for detection
In Figure 14, we show the minimum mixing ratio for detection achieved by the SC algorithm for each molecule. The analogous figure for the case of the MC algorithm is Figure 9. As shown, the SC algorithm achieves lower minimum mixing ratios of CH4 and O3, as compared with MC at lower SNRs. For instance, in the case of methane, the SC is capable of detecting the molecule at an SNR as low as 2, while the MC requires for the same mixing ratio a minimum SNR of 6. The case of ozone is also interesting, since incredibly low values of the mixing ratio of the molecule, namely $\sim 10^{-8}$, are identifiable at SNRgreater-$\gtrsim$18 with the SC algorithm.
Figure 14: Minimum mixing ratio required for detection of each molecule (with Recall $>$ 60%) with SC across different SNR values.
@@ 6.3.4 SC interference analysis
In Figure 15, we perform an analysis of the effect that interfering molecules have on the detection of each molecule for the SC algorithm. This analysis corresponds specifically to the case when the molecule of interest is absent and an interfering one can confuse the algorithm. A similar analysis was performed for the MC algorithm and presented in Figure 10. For details on the design of both figures, please refer to the explanation in subsubsection 6.2.4.
Figure 15: Same as Figure 10, but for the case of SC algorithms. For an animated version click here.
Compared to the MC case, the SC algorithm tends to be more conservative in correctly classifying the absence of all target molecules. In other words, SC tends to be less confused, especially for the case of O3 classification (middle panel in Figure 15). The same can be said for H2O classification (right-most panel), which, contrary to what is seen with the MC algorithm, is not confused by the presence of O3.
It is noteworthy that, unlike the MC case, relatively high SNR levels do not reduce the confusion of these molecules at high mixing ratios. This suggests that the SC takes advantage of the noise to avoid interference. However, this restricts the generalisation of the algorithm to other noise models, since the algorithm tends to depend on the noise structure and not on the signal itself for this particular task.
@@
Finally, in Figure 16, we analyze the effect of interference on the Recall metric, specifically for the positive detection of each molecule. The corresponding figure for the MC algorithm is presented in Figure 11.
Figure 16: Grid of Recall for SC at SNR=3. Same as Figure 11. For an animate version of this figure click here.
We notice that, in general, at low SNR, the behaviour of both algorithms is similar, at least for the biosignature molecules, ozone and methane. When the mixing ratio of these molecules is high, no interference from the other molecules is observed (upper and middle rows). At lower values of the mixing ratios, only in the case of methane, the presence of ozone can be helpful in positively identifying the molecule (upper row, left-most panel). This behaviour was also observed when using the MC algorithm.
However, for the classification of H2O using the SC algorithm (lower row), the presence of other molecules negatively affects the Recall. This is contrary to what happens when using the MC algorithm.
@@ 7 The SNR value and the number of transits
Throughout the paper, we have used SNR to quantify the precision of the transmission spectra when training and testing the algorithms. In a real-life scenario, for example, when performing observations using JWST, the level of noise of an observed spectrum will be determined by the number of transits that can be measured during an observing program or a survey campaign. In order to compare our results to those obtain in other works, we should estimate a relationship between the number of transits and the SNR of the resulting transmission spectra, at least for the scientific case we are studying here.
@@
In Figure 17 we show the correspondence between the number of transits of TRAPPIST-1 e, hypothetically observed with the JWST NIRSpec PRISM, and the resulting SNR as computed with Pandexo (Batalha et al., 2017). For making this plot, we first calculate, using TauREX, a theoretical spectrum for a TRAPPIST-1 e twin having a N2 atmosphere with CO2 at a mixing ratio of 10-2. Then, we synthesise the observed transmission spectrum of the planet using the observing program and instrument parameters chosen by Lin et al. (2021). We repeat the calculation assuming a number of transits between 1 and 100.
@@
Figure 17: SNR characteristic values (as computed at 4.28 $\mu$m) of the transmission spectra of a TRAPPIST-1 e twin with a N2-CO2 atmosphere as a function of the number of transits observed with the JWST NIRSpec PRISM. Horizontal lines indicate the detection thresholds for H2O (blue), CH4 (yellow), and O3 (red). Dashed lines represent the MC, while dotted lines represent the SC. The graph demonstrates that achieving an SNR above 10 would require an excessive number of transits, highlighting the challenges in detecting trace gases in exoplanetary atmospheres.
@@
Since the noise of this instrument depends on wavelength, defining a global SNR is not trivial. For our analyisis, we are interested on the peaks of the spectrum which provides information about the presence of molecules. It is more reasonable to use the SNR of the signal at some of those peaks than at other parts of the spectrum (for instance at the Rayleigh downhill). For this reason we use as a characteristic SNR, the value corresponding to the peak of CO2 at 4.28 $\mu$m. This peak is close enough to the maximum detectable wavelength of the instrument where the noise is also larger. We understand that, given the complexities of the instrumentation, the actual SNR value may change when different conditions are assumed.
@@
It is not hard to verify that the SNR value in Figure 17 grows as $\sqrt{N}$, where $N$ is the number of transits. This behaviour, although not self-evident in the figure since we have used a semi-log scale, it is the expected one since an increase in the number transits contribute to also increase the number of data points we add at a given channel in the measured flux. The main consequence of this behavior is that for increasing by a factor of 2 the value of the SNR we need almost 4 times more transits. This is one of the most important instrumental challenges we face when searching for biosignatures.
@@
If we are able develop, as proposed in this paper, a methodology able to select interesting targets having transmission spectra with SNR$<$3 for methane and SNR$<$24 for ozone, a significant improvement in our capabilities for searching for biosignatures could be achieved. This is precisely the ranges of SNR values where we have performed our numerical experiments.
To illustrate the capabilities of our method, we have included in Figure 17 the SNR value required to confidently classify planets at different biosignature mixing ratios, using the MC and SC algorithms. We see that for classifying a planet as interesting when containing levels as low as $10^{-7}$ of ozone, the MC algorithm requires almost 50 transits, while the SC needs only about 30 transits. Although a direct comparison is hard, and 30 transits seems to be a lot of them, it is interesting to notice that this number of transits is almost 4-6 times fewer than those required for standard retrieval procedures. On the other hand, and in the case of methane, the number of transits required for a reliable classification using our methods is similar to those achieved with some standard retrieval procedures. However, the mixing ratio reachable by our algorithms is typically 10 times smaller.
In summary, values of the SNR between 3 and 5, where our algorithms reach acceptable performances, correspond to a number of transits between 10 and 30 which are achievable for a JWST regular observing program. With a larger number of transits ($30-50$), our methods can classify planets as interesting at lower mixing ratios of methane than those required for regular retrieval procedures using the same number of transits. The case of ozone is more interesting, the MC and SC algorithms may tag a planet as interesting using as much as 6 times fewer transits than those required for normal retrieval procedures.
@@ 8 Classifying realistic spectra
In order to test our algorithms with realistic cases, we use the high-resolution spectra of the Earth at different geologic eras calculated by Kaltenegger et al. (2020). For those spectra we synthesise observations for the JWST NIRSpec using Pandexo. The synthesis followed the same procedure we apply to calculate the relationship between SNR value and number of transits in section 7.
@@
It is important to stress that the realistic spectra we use as inputs for these experiments, were originally calculated for a planet with a radius and surface gravity identical to that of the Earth. For using those spectra to study the slightly larger TRAPPIST-1 e planet, we simply rescale the line depth (expressed in kilometres) to the radius of our planet. Of course, the photochemical equilibrium composition of the Earth’s atmosphere and TRAPPIST-1 e, that orbits a M-dwarf star and will have also a different gravity and surface pressure, will be different. Still, since we we are just testing the performance of our algorithm when dealing with an actual planetary spectrum, the assumption of similar atmospheres are not violating any physical law.
@@
For a given number of transits (a given value of characteristic SNR), we synthesise 1000 observations and use the algorithms we trained in previous sections, namely the binary classificator (BC, subsection 6.1), the multilabel classificator (MC, subsection 6.2) and the specialised classificators (SC, subsection 6.3), to classify the planet as interesting or not interesting, or as having or not having specific molecules, respectively.
@@
In Figure 18, we show the value of the Recall metric, as obtained for the set of synthetic spectra of a TRAPPIST-1 e planet having a modern and a Proterozoic Earth’s atmosphere.
@@
Figure 18: Recall as a function of the number of transits calculated for a set of 1000 synthetic transmission spectra of a TRAPPIST-1 e twin, generated starting with the high-resolution spectra of the Earth’s modern atmosphere (upper panel) and Earth’s Proterozoic atmosphere (lower panel). Dotted lines with squares and dashed lines with triangles represent the performance of MC and SC classificators for CH4, O3, and H2O. The BC algorithm is shown as a continuous (purple) line.
@@
We observe that when dealing with more realistic spectra, namely an atmosphere with a more complex physical structure and a greater number of molecules, our binary classificator (BC) still performs well even if the number of transits is as low as the minimum of 1 transit. This promising result suggests that if we design a survey for obtaining low SNR transmission spectra, and use machine-assisted algorithms aimed at classifying the planets instead of attempting full retrieval as those proposed and exemplified in this paper, we could efficiently select planets that deserves follow-up observations. Interestingly, our results shows that even an algorithm that was trained only with synthetic and simple spectra, can correctly classify more than 60% of the planets with biosignatures in a survey sample. We speculate, that with a more complex training set and probably more competent methods, the performance of machine learning models using our methodology could be significantly enhanced.
Interestingly, the multilabel classificators (MC and SC) performs worse when dealing with realistic spectra, than the binary classificator (BC). For example, the MC classification of CH4, achieve acceptable performances (Recall $>0.6$) only at low number of transits, while the SC algorithm never reaches Recall values above the threshold, neither for the modern nor for the past Earth’s atmosphere. This result does not reveal a limitation of our general strategy but the fact that, in the case of multilabel classification, not training the algorithms for dealing with a realistic, instrument-specific noise, penalize the more these strategies. We may have identified the reason of this at least in the case of the SC algorithm. As we pointed out in subsubsection 6.3.4 the specialized algorithms performance may depend on the noise structure. Therefore, to be more efficient in an actual survey, SC algorithms should be properly trained, taking into account the noise of the instrument.
Finally, it is interesting to notice that, counter intuitively, the classification of O3 becomes harder as the number of transits increases. This fact suggests that the low levels of noise achieved in all spectral bands achieved after many transits, may strongly interfere with the classification of this molecule. Conversely, the classification performance for H2O and CH4 improves significantly with the number of transits, demonstrating the capabilities of the algorithms to handle transmission spectra with many lines.

### v4

@@ 7.3.1 SC multilabel metrics
In Figure 14, we show the result of calculating the multilabel classification metrics for the SC, using a common autoencoder (upper panel) and a specialised autoencoder strategy (lower panel). The latter strategy employs a set of autoencoders that are trained on spectra containing only one molecule at a time. This contrasts with the common autoencoder, used for our BC and MC algorithms, which is trained on spectra containing all types of molecules. Although the specialised autoencoder strategy is more CPU intensive—since, when classifying a spectrum with respect to a given molecule, the corresponding specialised autoencoder is required—there is a trade-off between a higher classification overhead and lower CPU usage during training with smaller datasets.
When comparing the results of the SC with those of the MC (see Figure 9), we note that despite differences in training set complexity and scalability, the performance of both approaches is nearly identical for most metrics. This implies that training a set of specialised classifiers, with the advantage of easily adding new molecules as needed, yields results comparable to those of a more complex and CPU-intensive algorithm that handles all molecules simultaneously (the MC algorithm).
[…]
In a more general case with a realistic spectrum, the MC approach will show its advantages over the SC, as we will see in section 9.
Figure 14: Multilabel metrics for the SC algorithm as a function of SNR. The top panel shows results obtained with the common autoencoder (AE), while the bottom panel corresponds to the specialised AE. The most notable difference between these approaches is seen in the chemical disequilibrium recall metric, which improves significantly under the specialised AE strategy. For an explanation of the shaded regions, see Figure 5.
@@ 7.3.2 SC molecule metrics
In Figure 15, we perform a similar comparison but evaluate the metrics for each molecule separately. This figure shows the results obtained with the common autoencoder strategy; the results for the specialised autoencoder strategy are not significantly different. The analogous results for the MC were presented in Figure 10.
Again, the performance of both algorithms is comparable, at least for methane and ozone. In the case of water, however, the SC algorithm performs poorly compared to the MC. This finding implies that, when classifying molecules individually, the more scalable SC algorithm can be as effective as the MC algorithm for most molecules.
Figure 15: Molecule-specific metrics from the SC using a common autoencoder. Each subplot shows the performance of different metrics (F1 Score, Recall, Precision, TNR) for each molecule (CH4, O3, H2O) as the SNR increases. For an explanation of the shaded regions, see Figure 5.
@@ 7.3.3 SC minimum mixing ratio for detection
In Figure 16, we show the minimum mixing ratio for detection achieved by the SC algorithm for each molecule. Here, we present the results obtained with the specialised autoencoder strategy, which outperforms the common autoencoder. The analogous figure for the MC algorithm is shown in Figure 11. As shown, the SC algorithm achieves slightly lower minimum mixing ratios for CH4 and O3 compared with the MC at lower SNRs.
Figure 16: Minimum mixing ratio required for the detection of each molecule (with Recall > 60%) using SC classifiers and specialised autoencoders across different SNR values.
@@ 7.3.4 SC interference analysis
In Figure 17, we analyse the effect that interfering molecules have on the detection of each target molecule for the SC algorithm. This analysis specifically addresses cases where the molecule of interest is absent and an interfering molecule may confuse the algorithm. A similar analysis was performed for the MC algorithm and is presented in Figure 12. For details on the design of both figures, please refer to the explanation in subsubsection 7.2.4.
Figure 17: Same as Figure 12, but for the case of SC algorithms with a common autoencoder.
Compared to the MC case, the SC algorithm tends to be more conservative at correctly classifying the absence of target molecules—that is, the TNR values are consistently above the threshold for all tested interfering mixing ratios.
Interestingly, and in contrast to the MC case, large amounts of interference from other molecules improve the algorithm’s performance in the detection of ozone and water (as evidenced by bumps at large mixing ratios in the middle and rightmost panels). This effect can be explained by considering that an algorithm specialised in detecting a given set of spectral signatures (for instance, those produced by ozone) will tend to recognise when significant spectral features of a different molecule are present, thus discounting those bands as belonging to the target molecule. It is as if the algorithm is "blinded" by the interfering molecule. This phenomenon is well illustrated by the differences observed in the interference on ozone detection produced by methane and water (middle panel). In particular, at a given mixing ratio, methane has a larger blinding effect on the ozone SC, likely because the spectral features of methane are more salient than those of water, at least in the studied spectral band. This is well illustrated in the spectrum shown in Figure 3.
@@
In Figure 18, we analyse the effect of interference on the Recall metric, specifically for the positive detection of each molecule. The corresponding figure for the MC algorithm is presented in Figure 13.
Figure 18: Grid of Recall for SC with specialized autoencoders at SNR=3. Same as Figure 13.
As expected from the behaviour of the TNR observed in Figure 17, the SC algorithms tend to respond negatively to the presence of interfering molecules (darker strips at high mixing ratios in the panels of the first column). However, this interference is less pronounced when the interfering molecule is water (panels of the second column).
The significant mutual interference between ozone and methane observed in the SC algorithm explains why the Chemical Disequilibrium Recall for this family of algorithms is lower than that of the MC (see Figure 14). In other words, in planets where both molecules are present, if one exists at a high mixing ratio, the SC algorithm tends to classify the planet as if the other molecule were absent, potentially discarding the most relevant astrobiological cases.
7.3.5 SC for airless and CO2-rich planets
Applying the same analysis as in subsubsection 7.2.5, but now for SC classifiers, we find that for airless planets both autoencoder approaches perform well, though notable differences exist. The specialised autoencoder yields a TNR near 0.9 for all molecules at SNR = 1, whereas the common autoencoder produces more variable results. At SNR = 3, stellar contamination can reduce the TNR for O3 and H2O to about 0.8 in the worst cases; beyond this point, performance improves as SNR increases. For CH4, the common autoencoder shows particularly strong performance, reaching TNR values above 0.9 at SNR = 5.
In CO2-dominated atmospheres, CH4 classification remains stable, with TNR around 0.7 at SNR = 1 for both autoencoder strategies and exceeding 0.8 at SNR = 5. O3 and H2O exhibit more strategy-dependent behaviour: at SNR = 1, the common autoencoder achieves a TNR of about 0.6 in the most contaminated cases, while the specialised approach maintains values above 0.65. Although TNR improves with increasing SNR for both strategies, the performance gap narrows from SNR = 3 onward.
@@ 8 The SNR value and the number of transits
Throughout the paper, we have used SNR to quantify the precision of the transmission spectra when training and testing the algorithms. In a real-life scenario, for example when performing observations using JWST, the noise level of an observed spectrum will be determined by the number of transits that can be measured during an observing program or survey campaign. In order to compare our results to those obtained in other works, we should estimate a relationship between the number of transits and the SNR of the resulting transmission spectra, at least for the scientific case we are studying here.
@@
In Figure 19 we show the correspondence between the number of transits for TRAPPIST-1 e, hypothetically observed with the JWST NIRSpec PRISM, and the resulting SNR as computed with PandExo (Batalha et al., 2017). To create this plot, we first calculate, using TauREx, a theoretical spectrum for a TRAPPIST-1 e twin having an N2 atmosphere with CO2 at a mixing ratio of $10^{-2}$. Then, we synthesise the observed transmission spectrum of the planet using the observing program and instrument parameters chosen by Lin et al. (2021). We repeat the calculation for a number of transits ranging from 1 to 100.
@@
Figure 19: Characteristic SNR values (computed at 4.28 $\mu$m) of the transmission spectra of a TRAPPIST-1 e twin with an N2-CO2 atmosphere as a function of the number of transits observed with the JWST NIRSpec PRISM. Horizontal lines indicate the detection thresholds for H2O (blue), CH4 (yellow), and O3 (red). The solid line represents the BC, while dashed lines correspond to the MC and dotted lines to the SC. This graph demonstrates that achieving an SNR above 10 would require an excessive number of transits, highlighting the challenges in detecting trace gases in exoplanetary atmospheres.
@@
Since the noise of this instrument depends on wavelength, defining a global SNR is not trivial. For our analysis, we are interested in the spectral peaks that provide information about the presence of molecules. It is more reasonable to use the SNR of the signal at one of these peaks than at other parts of the spectrum (for instance, at the Rayleigh slope). For this reason, we use as a characteristic SNR the value corresponding to the CO2 peak at 4.28 $\mu$m. This peak is sufficiently close to the maximum detectable wavelength of the instrument, where the noise is also higher. We acknowledge that, given the complexities of the instrumentation, the actual SNR value may vary under different conditions.
@@
It is not hard to verify that the SNR value in Figure 19 grows as $\sqrt{N}$, where $N$ is the number of transits. This behaviour—although not immediately evident in the figure due to the semi-log scale—is expected, as an increase in the number of transits contributes to an increase in the number of data points measured at a given channel in the flux. The main consequence of this behaviour is that doubling the SNR requires almost four times as many transits. This represents one of the most significant instrumental challenges we face when searching for biosignatures.
@@
If we are able to develop, as proposed in this paper, a methodology capable of selecting interesting targets with transmission spectra having SNR < 3 for methane and SNR < 24 for ozone, a significant improvement in our capabilities for searching for biosignatures/bioindicators could be achieved. These are precisely the SNR ranges at which we have conducted our numerical experiments.
To illustrate the capabilities of our method, we have included in Figure 19 the SNR required to confidently classify planets at different biosignature mixing ratios using the BC, MC and SC algorithms.
We find that, for classifying a planet as interesting when it contains ozone at levels as low as $10^{-7}$ — nearly one order of magnitude lower than present Earth levels, and the minimum detectable mixing ratio by our algorithms — the MC algorithm requires 30–40 transits, the BC algorithm 20–30, while the SC requires only 10–20 transits.
Although a direct comparison with other detection strategies is challenging, it is noteworthy that our best algorithm requires almost 6–8 times fewer transits for the reliable classification of most interesting planets than do standard retrieval procedures (see, e.g., Lin et al. 2021).
In the case of methane, the number of transits required for reliable classification using our methods (i.e., 10–20 transits) is similar to that achieved by some standard retrieval procedures. However, the minimum mixing ratio detectable by our algorithms is typically ten times lower. While this represents an extreme case, it illustrates the power of our approach.
In summary, SNR values between 3 and 5—where our algorithms reach acceptable performance—correspond to a number of transits between 10 and 30, which is feasible for a regular JWST observing programme. With a larger number of transits (30–50), our methods can classify planets as interesting at lower methane mixing ratios than those required by standard retrieval procedures using the same number of transits. The case of ozone is even more compelling; the MC and SC algorithms may tag a planet as interesting using up to six times fewer transits than those required for conventional retrieval procedures.
@@ 9 Classifying realistic spectra
In order to test our algorithms with realistic cases, we use the high-resolution spectra of Earth at different geologic eras calculated by Kaltenegger et al. (2020). For these spectra, we synthesise observations for JWST-NIRSpec using Pandexo. The synthesis followed the same procedure we apply to calculate the relationship between SNR and the number of transits in section 8.
@@
It is important to stress that the realistic spectra used as inputs for these experiments were originally calculated for a planet with a radius and surface gravity identical to Earth’s. To apply these spectra to the slightly larger TRAPPIST-1 e, we simply rescale the line depths in the original spectrum (expressed in kilometres) to the radius of our planet. Of course, the photochemical equilibrium composition of Earth’s atmosphere and that of TRAPPIST-1 e—which orbits an M-dwarf star and has different gravity and surface pressure—will be different. Nonetheless, since our purpose here is to test the performance of our algorithm when dealing with an actual planetary spectrum, assuming similar atmospheres does not violate any physical law.
For a given number of transits (i.e. a given characteristic SNR), we generate a sample of synthetic observations. To this end, we add stellar contamination to the theoretical signal using the procedure described in section 6. For each contaminated spectrum, we then synthesise 1,000 observed spectra, yielding a total of 10,000 test signals for each SNR (number of transits) value.
@@
Once we have the test samples, we use the algorithms trained in previous sections (see subsection 7.1, subsection 7.2 and subsection 7.3) to classify the planet as interesting or not interesting, or as having or not having specific molecules, respectively.
@@
In Figure 20, we show the Recall metric obtained for a set of synthetic spectra of a TRAPPIST-1 e planet with a modern Earth’s atmosphere (upper panel) and a Proterozoic Earth’s atmosphere (lower panel).
@@
Figure 20: Recall as a function of the number of transits calculated for a set of 10,000 synthetic transmission spectra of a TRAPPIST-1e twin, covering a wide range of stellar contamination scenarios. These spectra were generated starting from high-resolution spectra of Earth’s modern atmosphere (upper panel) and Earth’s Proterozoic atmosphere (lower panel). Dotted lines with squares and dashed lines with triangles represent the performance of MC and SC classifiers for CH4, O3, H2O, and chemical disequilibrium (simultaneous CH4+O3). The BC algorithm is shown as a continuous (purple) line.
In previous analysis we showed the value of the metrics obtained for different degrees of contamination considered separately. This gave us the shaded bands in Figure 7-Figure 17. For this analysis, and for the sake of simplifying the analysis of the results, we obtain the recall of the full set, independently of the level of contamination.
We observe that when dealing with more realistic spectra, namely an atmosphere with a more complex physical structure and a greater number of molecules, the binary (BC) and multilabel classifiers (MC) perform well even if the number of transits is as low as 4-10.
This result suggests that small-to-medium-scale observing campaigns, combined with classification-oriented machine-learning algorithms rather than full atmospheric retrievals, can effectively identify planets warranting detailed follow-up observations.
@@
Our results demonstrate that even an algorithm that was trained only with synthetic and simple spectra, can correctly classify more than 60% of the planets with biosignatures/bioindicators in a survey sample. We speculate, that with a more complex training set and probably more competent methods, the performance of machine learning models using our methodology could be significantly enhanced.
The specialized classifier (SC, dotted lines) performs worse when dealing with realistic spectra, than the binary (BC, continuous line) and multilabel (MC, dashed lines) classifier.
This outcome arises because our specialised algorithm is overfitted to classify less complex spectra with a simplified noise structure,. However, the realistic spectra we are using for this test has many molecular species and a wavelength-dependent instrumental noise. It is interesting, however, to notice how well the BC and MC algorithms performs when dealing with such complex spectra, even although they were also trained with less realistic signals.
As expected, the values of the SNR where the BC and MC algorithms reach the threshold of 60% spectra correctly classified, are consistent with the minimum mixing ratio for detection (see Figure 6 and Figure 11) for the levels of methane and ozone in the modeled Earth.